how do i solve
2x-5y=-12 and
-4x+y=6
using substitution?

Answer by elima(1433)

First solve one of the equations for one of the variables;
-4x+y=6
y=6+4x
Now substitute that into the other equation and solve for x;
2x-5y=-12
2x-5(6+4x)=-12
2x-30-20x=-12
-30-18x=12
-18x=-12+30
-18x=18
x=-1
Now plug that into the equation to find y;
-4x+y=6
-4(-1)+y=6
4+y=6
y=6-4
y=2
Hope you understand
=)